### v1

# RIGNO: A Graph-based framework for robust and accurate operator learning for PDEs on arbitrary domains

Abstract
Learning the solution operators of PDEs on arbitrary domains is challenging due to the diversity of possible domain shapes, in addition to the often intricate underlying physics. We propose an end-to-end graph neural network (GNN) based neural operator to learn PDE solution operators from data on point clouds in arbitrary domains. Our multi-scale model maps data between input/output point clouds by passing it through a downsampled regional mesh. Many novel elements are also incorporated to ensure resolution invariance and temporal continuity. Our model, termed RIGNO, is tested on a challenging suite of benchmarks, composed of various time-dependent and steady PDEs defined on a diverse set of domains. We demonstrate that RIGNO is significantly more accurate than neural operator baselines and robustly generalizes to unseen spatial resolutions and time instances.
1 Introduction
Partial Differential Equations (PDEs) mathematically model a wide variety of interesting phenomena in physics and engineering, making their study a matter of great scientific importance [9]. Given the paucity of explicit solution formulas, numerical methods [34], such as finite difference, finite element and spectral methods constitute the main tools in the quantitative study of PDEs. Unfortunately, the computational cost of existing numerical methods or solvers for PDEs can be prohibitively expensive, particularly in the context of problems such as uncertainty quantification, inverse problems, control, and design, where PDE solvers have to be called multiple times [35]. This necessitates the design of fast and accurate surrogates for PDE solvers.
In this context, machine learning (ML) and AI-inspired algorithms are increasingly being used as surrogates for PDE solvers, see [30] and the references therein. In particular, operator learning, i.e., where the task at hand is to learn the underlying solution operator of a PDE from data, has emerged as the dominant paradigm in the application of ML/AI techniques for PDEs. A large variety of operator learning algorithms have recently been proposed, with examples being DeepONets [7, 27], Fourier Neural Operator (FNO) [20], Low-rank Neural Operators [19], Convolutional Neural Operators (CNO) [37] and transformer-based operator learning algorithms [6, 33, 10, 12, 1], among many others.
An overwhelming majority of these operator learning algorithms or neural operators [15] assume that the underlying domain is Cartesian and is discretized on a regular grid; however, in practice, and particularly in engineering, PDEs are posed on domains with very complex and highly non-Cartesian geometries. Thus, the direct application of many of the aforementioned operator learning algorithms is difficult or impractical. Therefore, we require algorithms that can learn the solution operators of PDEs on arbitrary domains.
Various approaches have been proposed to address this question. To start, methods such as the FNO and the CNO, originally designed for Cartesian domains, can be extended to non-Cartesian domains via domain masking [37, 12]. Another approach for extending FNO to arbitrary domains is to directly evaluate the discrete Fourier transform with batched matrix multiplications rather than FFT [23]. Moreover, one can transform the underlying problem on a non-Cartesian domain to one on Cartesian domains either through extensions [14], learned diffeomorphisms as in Geo-FNO [21], or through learned interpolations between non-Cartesian physical domain and a Cartesian latent domain via graph neural operators [18] as in GINO [22] and references therein.
A different approach to learning operators for PDEs on arbitrary domains is provided by graph neural networks or GNNs [5]. This is an ideal ML framework for handling relational data such as molecules, networks, or unstructured computational grids. The use of graph neural networks for simulating PDEs has received widespread attention [38, 39, 32, 4, 8], to name a few. However, genuine learning of operators, rather than a specific finite-dimensional discretization of them, requires the ability of the underlying algorithm to display some form of resolution invariance [15, 2], i.e., the operator learning algorithm should be able to process inputs and outputs on different spatial resolutions, both during training and testing, without any decrease in its overall accuracy. A priori, it is unclear if the existing GNNs for simulating PDEs on arbitrary domains satisfy this key property of resolution invariance.
Another key aspect of operator learning for PDEs is the issue of temporal discretization when approximating solution operators of time-dependent PDEs. These operators are continuous in time and the algorithm approximating them should be able to retain this continuity. However, existing frameworks for learning time-dependent PDEs are often based on auto-regressive rollouts [4, 38, 39, 8] and cannot be evaluated at arbitrary times without additional interpolations, limiting their applicability in the context of time-dependent PDEs.
Thus, there is a pressing need for an ML algorithm that can learn solution operators of PDEs on arbitrary domains, which can also deal with multiple spatial resolutions and provide appropriate temporal behavior when used for time-dependent PDEs, while simultaneously providing precise and accurate approximations to the underlying operator. The design of such an algorithm is the main point of this article.
To this end, we propose Region Interaction Graph Neural Operator (RIGNO) as an effective neural operator to learn PDEs on arbitrary domains. RIGNO uses GNNs end-to-end and does not require any transformations into Cartesian domains at any stage. It is constructed for multiscale information processing and many novel elements are introduced to enable RIGNO to provide spatial resolution invariance and to directly and continuously approximate the solution operator of time-dependent PDEs for any time. Through extensive numerical experiments, we demonstrate that RIGNO provides extremely accurate approximations to the solution operators of a variety of PDEs with domain geometries including Cartesian grids, structured and unstructred meshes, and random point clouds. Thus, we present a GNN-based framework for accurately and robustly learning the solution operators of PDEs on arbitrary domains.
2 Methods
Problem Formulation. We consider the following generic time-dependent PDE with $\Omega_{t}=(0,T)$ the time domain, $\Omega_{x}\subset\mathbb{R}^{d}$ the $d$-dimensional spatial domain, $u\in\mathcal{C}(\overline{\Omega}_{t};\mathcal{X})$ the solution of the PDE with $\mathcal{X}\subset L^{p}(\Omega_{x};\mathbb{R}^{s})$ for some $1\leq p<\infty$, $c\in\mathcal{Q}$ a spatially varying coefficient, $\mathcal{F}$ a differential operator, $\mathcal{B}$ a boundary operator, and $a\in\mathcal{X}$ the initial condition,
$\partial_{t}u=\mathcal{F}(c,t,u,\nabla_{x}u,\nabla^{2}_{x}u,\dots),\quad\forall(t,x)\in\Omega_{t}\times\Omega_{x},$ (1)
$\mathcal{B}u(t,x)=0,\quad\forall(t,x)\in\Omega_{t}\times\partial\Omega_{x},$
$u(0,x)=a(x),\quad\forall x\in\Omega_{x}.$
The solutions of the PDE (1) are given in terms of the solution operator $\mathcal{S}^{t}:\mathcal{X}\times\mathcal{Q}\times\Omega_{t}\to\mathcal{X}$, which maps the initial conditions and coefficients to the solution $u(t)$ at any time. With the semi-group property for solutions of time-dependent PDEs [9], we can define an extended solution operator $\mathcal{G}^{\dagger}:\mathcal{X}\times\mathcal{Q}\times\Omega_{t}\times\mathbb{R}^{+}\to\mathcal{X}$, which maps the solution $u(t)$ at any time $t$ and the coefficient $c$ to the solution $u(t+\tau)$ at a later time $t+\tau$ through the action,
$\mathcal{G}^{\dagger}(u(t),c,t,\tau)=\mathcal{S}^{\tau}(u(t),c,\tau)=u(t+\tau).$ (2)
Our goal here is to learn this extended solution operator $\mathcal{G}^{\dagger}$ from data. Note that plugging $t=0,\tau=t,u(0)=a$ yields $\mathcal{S}^{t}$, recovering the solution operator of the PDE (1). We would also like to point that the (extended) solution operator depends implicitly on the underlying spatial domain $\Omega_{x}$, which in turn, can change in some operator learning tasks.
Model Architecture.
As shown in Figure 1, our model is based on the well-known encode-process-decode paradigm [38, 39]. All the three components, i.e., encoder, processor, and decoder are graph neural networks (GNNs). The methods presented in this section are implemented using JAX [3] and are available on Github: https://github.com/camlab-ethz/rigno. As the aim is to learn the (extended) solution operator $\mathcal{G}^{\dagger}$ (2), the inputs to our model are, in general, i) a set of coordinates in the domain $\boldsymbol{\rm x}:=\{\boldsymbol{\rm x}_{i}\in\Omega_{x}\}_{i}$, where $\Omega_{x}$ is the spatial domain, ii) $\boldsymbol{u}:=\{\boldsymbol{u}_{i}\in\mathbb{R}^{s}\}_{i}$, discretized values of a function at known coordinates $\boldsymbol{\rm x}$, iii ) $\boldsymbol{c}:=\{\boldsymbol{c}_{i}\}_{i}$, discretized values of the spatial coefficient at $\boldsymbol{\rm x}$ coordinates, iv) the time $t$ and v) the lead time $\tau$. The model’s output $\boldsymbol{r}:=\{\boldsymbol{r}_{i}\in\mathbb{R}^{s}\}_{i}$ represent discretized values of a parameterized map at the coordinates $\boldsymbol{\rm x}$,
$\boldsymbol{r}=\mathcal{T}_{\theta}(\boldsymbol{u},\boldsymbol{\rm x},\boldsymbol{c},t,\tau).$ (3)
Underlying Graphs.
The map $\mathcal{T}_{\theta}$ is a GNN and will process data that are defined on graphs. As shown in Figure 1 (for an idealized two-dimensional domain) and following ideas in [16], we consider two underlying sets of nodes. The first set of nodes corresponds to the unstructured point cloud on which the input (and output) data is defined. We call this point cloud as the physical mesh and the nodes that constitute it as physical nodes. We denote them as $\boldsymbol{V}_{\mathfrak{P}}=\{\boldsymbol{v}_{i}^{\mathfrak{P}}\}_{i}$. Note that we do not require any explicit relational information on these physical nodes.
Most of our processing will take place in another bespoken set of nodes that we call as regional nodes and the point cloud constituted by them as the regional mesh, denoted as $\boldsymbol{V}_{\mathfrak{R}}=\{\boldsymbol{v}_{i}^{\mathfrak{R}}\}_{i}$. Here, the regional nodes are constructed by randomly sub-sampling the set of physical nodes with a sampling ratio which is a hyperparameter of our model. A graph is constructed to connect the regional nodes (see Figure 1) and the resulting bidirectional regional mesh edges are denoted by $\boldsymbol{E}_{\mathfrak{R}\leftrightarrow\mathfrak{R}}=\{\boldsymbol{e}_{ij}^{\mathfrak{R}\leftrightarrow\mathfrak{R}}\}_{ij}$.
Moreover, we need to pass information between the physical nodes and the regional nodes, necessitating the construction of graphs between them. To this end, we connect physical nodes to regional nodes (Figure 1) by unidirectional edges $\boldsymbol{E}_{\mathfrak{P}\to\mathfrak{R}}=\{\boldsymbol{e}_{ij}^{\mathfrak{P}\to\mathfrak{R}}\}_{ij}$. These edges are constructed such that each regional node constitutes a sub-region for a subset of (nearby) physical nodes and serves to accumulate information from them. Similarly, information from the regional mesh can be transmitted to the physical mesh by constructing a graph with unidirectional edges $\boldsymbol{E}_{\mathfrak{R}\to\mathfrak{P}}=\{\boldsymbol{e}_{ij}^{\mathfrak{R}\to\mathfrak{P}}\}_{ij}$. As shown in Figure 1, these edges are constructed to accumulate information from multiple regional nodes into one physical node. Thus, construction of these unidirectional edges provides a very flexible upscaling and downscaling strategy for our model.
Figure 1: General schematic of the RIGNO architecture for an idealized two-dimensional domain. The inputs are first independently projected to a latent space by feed-forward blocks. The information on the original discretization (physical nodes) is then locally aggregated to a coarser discretization (regional nodes). Regional nodes are connected to each other by edges with multiple length scales. Several message-passing steps are then applied on the regional nodes which constitute the processor. The processed features are then transmitted back to the original discretization by using similar edges as in the encoder, before being independently projected back to the desired output dimension via a feed-forward block without normalization layers.
Needless to say, when possible, the graph construction must be adapted to the boundary conditions of the PDE (1). We postpone the rather technical description of graph construction and incorporation of boundary conditions, particularly the novel additions to the model design to enforce periodic boundary conditions, to SM Section A.
Once the underlying graphs have been constructed, we must also provide geometric information to the GNNs. To this end, we impose translation equivariance by employing linear normalized coordinates $\boldsymbol{\zeta}_{i}\in[-1,+1]$ and angular normalized coordinates $\boldsymbol{\alpha}_{i}\in[0,2\pi]$ defined as
$\boldsymbol{\zeta}_{i}:=2\frac{\boldsymbol{\rm x}_{i}-\boldsymbol{\rm x}_{\rm c}}{\boldsymbol{L}}-1;\qquad\boldsymbol{\alpha}_{i}:=\pi(\boldsymbol{\zeta}_{i}+1).$
The structural features of an edge $\boldsymbol{e}_{ij}$ connecting node $i$ to node $j$ is $\boldsymbol{\varphi}_{ij}:=\left[\tilde{\boldsymbol{\zeta}}_{ij}\diamond\|\tilde{\boldsymbol{\zeta}}_{ij}\|_{2}\right]$; with $\tilde{\boldsymbol{\zeta}}_{ij}:=(\boldsymbol{\zeta}_{j}-\boldsymbol{\zeta}_{i})/(2\sqrt{d})$ being the relative coordinates of the connected nodes normalized by the diameter of the domain, $2\sqrt{d}$. We denote the vertical concatenation of column vectors by $\diamond$. The structural features of a node $\boldsymbol{v}_{i}$ with general boundary conditions are simply its normalized coordinates $\boldsymbol{\varphi}_{i}:=\boldsymbol{\zeta}_{i}$. With periodic boundary conditions (see SM Section B), we concatenate sines and cosines of angular normalized coordinates with $K_{\rm freq}$ frequencies:
$\boldsymbol{\varphi}_{i}=\left[\sin(k\boldsymbol{\alpha}_{i})\diamond\cos(k\boldsymbol{\alpha}_{i})\right],\quad 1\leq k\leq K_{\rm freq}.$ (4)
For regional nodes, the coverage radius of each node, $R_{i}$, is also concatenated with the other structural features. Once the inputs, graphs and geometric information are available, our model processes this information by employing the following constituent GNNs.
Encoder. Our encoder (Figure 1) is designed to transmit information from the underlying physical mesh (representing a discretization of the domain $\Omega_{x}$) to the (coarser) regional mesh, while at the same time, also embedding inputs into a latent space to enable representation learning. These are coupled into the following encoder module, consisting of two steps. In the first step, all the features are mapped into suitable latent spaces with,
[…]
Here and everywhere in this paper, FF denotes a feed-forward block (see Figure 1), which consists of a shallow multi-layer perceptron (MLP) followed by normalization layers. Next, the second step is a single message-passing block, with residual updates, to move the information from the physical mesh to the regional mesh:
$\boldsymbol{m}^{\mathfrak{P}\to\mathfrak{R},(0)}_{ij}=\textsl{FF}_{\text{encoder}}^{\mathfrak{P}\to\mathfrak{R}}(\boldsymbol{v}^{\mathfrak{P},(0)}_{i},\boldsymbol{v}^{\mathfrak{R},(0)}_{j},\boldsymbol{e}^{\mathfrak{P}\to\mathfrak{R},(0)}_{ij}),$ (6)
$\boldsymbol{v}^{\mathfrak{R},(1)}_{i}=\boldsymbol{v}^{\mathfrak{R},(0)}_{i}+\textsl{FF}_{\text{encoder}}^{\mathfrak{R}}({\boldsymbol{v}^{\mathfrak{R},(0)}_{i}},\tfrac{1}{|\mathcal{N}_{i}|}\sum\limits_{j\in\mathcal{N}_{i}}\boldsymbol{m}^{\mathfrak{P}\to\mathfrak{R},(0)}_{ji}),$
$\boldsymbol{v}^{\mathfrak{P},(1)}_{i}=\boldsymbol{v}^{\mathfrak{P},(0)}_{i}+\textsl{FF}_{\text{encoder}}^{\mathfrak{P}}({\boldsymbol{v}^{\mathfrak{P},(0)}_{i}}).$
Processor. The processor is designed to process the latent information on the regional mesh by $P$ sequential message-passing steps with residual updates and skip connections to allow for gradients to flow for large $P$, in order to facilitate stable training. The $p$-th message-passing step in the processor updates the regional nodes and the regional mesh edges as:
$\boldsymbol{m}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p-1)}_{ij}=\textsl{FF}_{\text{processor}}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p)}(\boldsymbol{v}^{\mathfrak{R},(p)}_{i},\boldsymbol{v}^{\mathfrak{R},(p)}_{j},\boldsymbol{e}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p-1)}_{ij}),$ (7)
$\boldsymbol{v}^{\mathfrak{R},(p+1)}_{i}=\boldsymbol{v}^{\mathfrak{R},(p)}_{i}+\textsl{FF}_{\text{processor}}^{\mathfrak{R},(p)}({\boldsymbol{v}^{\mathfrak{R},(p)}_{i}},\tfrac{1}{|\mathcal{N}_{i}|}\sum\limits_{j\in\mathcal{N}_{i}}\boldsymbol{m}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p-1)}_{ji}),$
${\boldsymbol{e}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p)}_{ij}}={\boldsymbol{e}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p-1)}_{ij}}+\boldsymbol{m}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(p-1)}_{ij}.$
Decoder. The decoder serves to transfer processed information from the regional mesh back to the physical mesh (Figure 1). To this end, the first step in the decoder is a single message-passing block:
$\boldsymbol{m}^{\mathfrak{R}\to\mathfrak{P},(0)}_{ij}=\textsl{FF}_{\text{decoder}}^{\mathfrak{R}\to\mathfrak{P}}(\boldsymbol{v}^{\mathfrak{R},(P_{\mathfrak{R}}+1)}_{i},\boldsymbol{v}^{\mathfrak{P},(1)}_{j},\boldsymbol{e}^{\mathfrak{R}\to\mathfrak{P},(0)}_{ij}),$ (8)
$\boldsymbol{v}^{\mathfrak{P},(2)}_{i}=\boldsymbol{v}^{\mathfrak{P},(1)}_{i}+\textsl{FF}_{\text{decoder}}^{\mathfrak{P}}({\boldsymbol{v}^{\mathfrak{P},(1)}_{i}},\tfrac{1}{|\mathcal{N}_{i}|}\sum\limits_{j\in\mathcal{N}_{i}}\boldsymbol{m}^{\mathfrak{R}\to\mathfrak{P},(0)}_{ji}).$
Finally, in the second step of the decoder, the physical node feature vectors are independently scaled down to the desired dimension to form the output:
$\boldsymbol{r}_{i}=\textsl{FF}_{\text{decoder}}^{\text{output}}(\boldsymbol{v}^{\mathfrak{P},(2)}_{i}).\$ (9)
Summarizing, as our model involves the learned interaction between nodes that each represent aggregated information from an underlying sub-region, we call it as Region Interaction Graph Neural Operator or RIGNO.
Edge Masking.
Inspired by dropout [40], we employ random masking of edges in each message-passing step of RIGNO. Before each message-passing step (e.g., step $p$ of the processor (7)), each edge is masked with a given probability and does not participate in that specific message-passing block. In the next message-passing step (e.g., step $p+1$ of the processor), a new set of edges is randomly masked. Since the messages computed from the neighbors are averaged, there is no need for any further configuration to apply this at training or inference.
As shown in the next section, edge masking enables us to i) make RIGNO (approximately) resolution invariant, ii) reduce computational complexity by sparsely sampling edges during each step, and iii) provide a method for quantifying uncertainty in the predictions of RIGNO.
Temporal Processing.
We recall that our goal is to approximate the extended solution operator $\mathcal{G}^{\dagger}(u(t),c,t,\tau)$ (2), which is a function of the input time $t$ and the lead time $\tau$. Both of these scalars are provided to the encoder of RIGNO as inputs and embedded into suitable high-dimensional latent spaces. However, merely providing lead time as an input may not be sufficient to learn the solution operator in deep models. Hence, we introduce lead time as an additional conditioning in each hidden layer of RIGNO (see Figure 1). We do this by adapting the normalization layers employed within our model. To this end, let ${\bf y}$ be an output vector for any layer and normalize it by an appropriate normalization layer (BatchNorm, LayerNorm, InstanceNorm, etc.) to obtain the normalized output $\widetilde{{\bf y}}$. We further condition it with the lead time $\tau$ by setting,
$\hat{{\bf y}}=(1+\tau\gamma(\tau))\odot\widetilde{{\bf y}}+\tau\lambda(\tau),$ (10)
with $\gamma,\lambda$ being small learnable MLPs.
Finally, we need to provide a time stepping strategy to update our model in time. In the context, RIGNO can be employed in a very general time stepping mode by setting,
$\mathcal{G}_{\theta}(u^{t},c^{t},t,\tau)=\alpha u^{t}+\beta\mathcal{T}_{\theta}(u^{t},c^{t},t,\tau),$ (11)
where $\alpha,\beta$ are scalar parameters, $\mathcal{T}_{\theta}$ is the output of RIGNO, and $\mathcal{G}_{\theta}\approx\mathcal{G}^{\dagger}$ is the final approximation to the extended solution operator (2) of the PDE (1).
Setting $\alpha=0$ and $\beta=1$ directly recovers an approximation to the output of the extended solution operator. In this case, RIGNO can be interpreted to directly provide the solution of the PDE (1) at time $t+\tau$, given the solution at time $t$, coefficient $c$ and the lead time $\tau$ as inputs. Similarly, setting $\alpha=\beta=1$ yields the residual of the solution at the later time, given the solution at current time. On the other hand, setting $\alpha=1$ and $\beta=\tau$ leads to the model output approximating the time derivative of the solution operator as (11) is simply an approximation of the Taylor expansion in time of the solution operator, truncated at leading order. A clear advantage of this time-derivative update is that it automatically enforces time continuity by simply setting $\tau=0$. We will explore these choices for time updates in our numerical experiments to determine which ones are the most suitable.
Loss Functions and Training Strategy.
For simplicity of exposition here, we fix a final time $T$ and a fixed discretization $\Omega_{t}=\cup_{n=1}^{N}[t_{n-1},t_{n}]$. We assume that training samples are provided as trajectories $\{u_{m}(t_{n})\}_{n=0}^{N}$, for $1\leq m\leq M$, with each $u_{m}$ being the solution to the PDE (1) with initial condition $a_{m}$ and coefficient $c_{m}$. Given these training trajectories and the fact that our model can be evaluated at any time, there are multiple ways in which the RIGNO model, denoted here by $\mathcal{G}_{\theta}$ can be trained. We follow the recently proposed all2all training strategy [12] as training input-output pairs scale quadratically with respect to the number of time steps for time-dependent problems. Hence, the goal of training is to find parameters $\theta\in\Theta$ such that the following loss function can be be minimized,
${\mathcal{J}}:=\frac{1}{M\widehat{N}}\sum_{m,n,t_{k}\leq t_{n}}\mathcal{L}\left[u_{m}({t_{n}}),\;\widehat{u}_{m}^{t_{k}}({t_{n}})\right],$ (12)
$\widehat{u}_{m}^{t_{k}}({t_{n}})=\mathcal{G}_{\theta}\left(u_{m}({t_{k}}),c_{m},\;t_{k},\;t_{n}-t_{k})\right),$
with $\widehat{N}=\frac{1}{2}(N+1)(N+2)$ and $\mathcal{L}$ being any measure of the mismatch between the ground truth and the model prediction, for instance either the mean absolute or the mean squared error.
A problem with this approach is that the smallest prediction lead time $\tau_{min}$ in the training set is given by the minimum of $t_{n}-t_{n-1}$, for all $n$. Hence, at inference, predicting for lead times $\tau<\tau_{min}$ could lead to distribution shifts as smaller lead times have not been seen during training. Yet, a fundamental requirement for a time-dependent neural operator is to predict arbitrarily small times correctly in order to ensure time continuity. We tackle this problem by proposing to further fine-tune a RIGNO model, trained with the loss function (12), by using a modified loss function where $\widehat{u}_{m}^{t_{k}}({t_{n}})$ in (12) is obtained by starting from an intermediate input function: $\mathcal{G}_{\theta}(\widetilde{u}_{m}({t_{\ast}}),c_{m},t_{\ast},t_{n}-t_{\ast}$). Thus, our aim is to use solutions at intermediate time steps, which are not available in the training set, to make predictions at later time steps, available in the training data. However, we must find an approximation $\widetilde{u}_{m}({t_{\ast}})\approx u_{m}({t_{\ast}})$ as $u_{m}(t_{\ast})$ is not available during training. Our solution is to use the pretrained RIGNO itself by setting,
$\tilde{u}({t_{\star}})=\mathcal{G}_{\theta}\left(u(t_{k}),\;c,\;t_{k},\;t_{\star}-t_{k}\right).$
This self-consistent approach is reasonable as the modified loss function is only used during fine-tuning. Moreover, as shown in SM Figure C.3, the flow of the gradient can be stopped at the intermediate time $t_{\star}$, similar to ideas introduced in [4], which will be referred to as GradCut. The utility of this fractional pairing strategy will be tested subsequently.
Given that the output time $t$ and lead time $\tau$ are both inputs to the trained RIGNO, a variety of approaches can be used for temporal processing at inference. As in [12], to obtain the solution $u(t)$ at time $t$, from an initial datum $u(0)=a$, we can either directly apply RIGNO with lead time $\tau=t$ or we can employ different autoregressive rollout strategies as described in the SM Section C.2.
Table 1: Benchmarks on datasets with unstructured and uniform-grid space discretizations. Lowest (blue) and second lowest (orange) errors are highlighted with colors. The errors for time-dependent problems are calculated with three different autoregressive rollouts; and the lowest error is reported for each architecture. The details of autoregressive rollouts are available in SM Section C.2. All trainings are done with 1024 samples, except for the AF dataset where 2048 training samples are used. We observed that the relative errors were very sensitive to the normalization applied on the target values. More details about the metric used are available in SM Section E.
Dataset Median relative $L^{1}$ error [%]
(unstructured) RIGNO-18 RIGNO-12 GeoFNO FNO DSE GINO
Heat-L-Sines 0.04 0.05 0.15 0.53 0.19
Wave-C-Sines 5.35 6.25 13.1 5.52 5.82
NS-Gauss 2.29 3.80 41.1 38.4 13.1
NS-PwC 1.58 2.03 26.0 56.7 5.85
NS-SL 1.28 1.91 24.3 29.6 4.48
NS-SVS 0.56 0.73 9.75 26.0 1.19
CE-Gauss 6.90 7.44 42.1 30.8 25.1
CE-RP 3.98 4.92 18.4 27.7 12.3
ACE 0.01 0.01 1.09 1.29 3.33
Wave-Layer 6.77 9.01 11.1 28.3 19.2
AF 1.00 1.09 4.48 1.99 2.00
Elasticity 4.31 4.63 5.53 4.81 4.38
(uniform grid) RIGNO-18 RIGNO-12 CNO scOT FNO
NS-Gauss 2.74 3.78 10.9 2.92 14.24
NS-PwC 1.12 1.82 5.03 7.11 11.24
NS-SL 1.13 1.82 2.12 2.49 2.08
NS-SVS 0.56 0.75 0.70 0.99 6.21
CE-Gauss 5.47 7.56 22.0 9.44 28.69
CE-RP 3.49 4.43 18.4 9.74 31.19
ACE 0.01 0.01 0.28 0.21 0.60
Wave-Layer 6.75 8.97 8.28 13.44 28.01
3 Results
In this section, we demonstrate the performance of RIGNO on a challenging suite of numerical experiments by comparing it with well-established baselines. We also highlight the factors that underpin the observed performance of RIGNO.
RIGNO accurately learns the solution operators of PDEs on arbitrary domains.
To test the performance of RIGNO for PDEs on arbitrary domains, we choose two types of datasets. First, we consider 4 datasets on structured and unstructured meshes, namely Heat-L-Sines (heat equation on a L-shaped domain), Wave-C-Sines (wave equation, propagation of waves in a homogeneous medium of circular shape), AF (compressible transonic flow past airfoils taken from [21]), and Elasticity (hyper-elastic deformations of varying domain shapes, also taken from [21]). See SM Section D for details of datasets and their abbreviations used here. Next, we test how well RIGNO learns operators defined on arbitrary point clouds. To set this up, we choose 8 datasets from the PDEGym database [12], which deal with four different PDEs; the incompressible Navier-Stokes equations, the compressible Euler equations, the Allen-Cahn equation, and the wave equation on a spatially varying medium. The datasets are all originally defined on uniform Cartesian grids. For the unstructured experiments, we randomly subsample 9216 points from the available coordinates on a 128$\times$128 uniform grid, see SM Figure D.1.
In Table 1, we present the test errors on these 12 unstructured datasets [31] with two sizes of RIGNO, namely RIGNO-12 and RIGNO-18, with 12 and 18 layers in their processor modules, respectively. To contextualize the results, we consider three baselines, namely Geo-FNO [21], FNO DSE [23] and GINO [22]. All of the baselines are capable of handling inputs and outputs on arbitrary point clouds, yet use different routes to achieve this. More details about the baselines, including the selection of hyperparameters, are outlined in SM Section F. We observe from Table 1 that RIGNO-18 outperforms all the baselines in all 12 datasets. Likewise, RIGNO-12 is the second best-performing model for 10 of the 12 datasets. For Wave-C-Sines and Elasticity, FNO DSE and GINO respectively claim second best, yet RIGNO-12 remains comparable in performance to RIGNO-18. Moreover, in some cases, the RIGNO variants are almost an order of magnitude more accurate than the baselines, showing excellent performance on these benchmarks. We would also like to emphasize that RIGNO is the only framework that displays acceptably low errors for every single dataset that we consider here, demonstrating its robustness.
Often, models are designed to have flexible inputs/outputs by potentially reducing their performance on Cartesian domains. To further illustrate the expressive power of RIGNO, we consider the same 8 datasets from the PDEGym database as before, but on their original Cartesian grids. In this case, we can compare RIGNO with state-of-the-art neural operators such as FNO, CNO, and scOT which are tailored for Cartesian grids. The resulting test errors, presented in Table 1, also show that RIGNO is very accurate on Cartesian grids and RIGNO-18 is the best performing model on every single dataset. Again, RIGNO-12 is either the second best-performing model or comparable to it. In fact, for most datasets, RIGNO-18 is several times more accurate than neural operators such as CNO and scOT, as well as the widely used FNO, showcasing the expressive power of its GNN-based backbone.
(a) Training dataset size.
(b) Model size.
Figure 2: Autoregressive test errors at $t=t_{14}$. Fractional pairing is off in all trainings. All results correspond to the AR-2 rollout scheme (defined in SM Section C.2). All axes are in logarithmic scales. The plot on the right corresponds to the NS-PwC dataset.
RIGNO scales with data and model size.
An essential requirement for modern ML models is their ability to scale with data, i.e., the test error should reduce at a consistent rate as the number of training samples increases [13, 17, 12]. To test this aspect, we plot the test error with respect to the number of training samples in Figure 2(a), observing that RIGNO-12 shows a consistent scaling for the datasets. Another critical aspect of ML models is their ability to scale with model size [13, 12]. Compared to data scaling, scaling with model size is more subtle as a number of hyperparameters influence model size and changing each of them, while keeping the others fixed, might lead to different outcomes. In SM Section G.2, we present a detailed discussion and results on how RIGNO scales with respect to each of the hyperparameters that influences model size. Here, in Figure 2(b), we only consider two (correlated) indications of model size, i.e., total number of model parameters and inference time. We see from Figure 2(b) that RIGNO-12 consistently scales with respect to both the total number of parameters and inference run time.
(a) Edge masking (NS-PwC).
(b) fractional pairing (NS-PwC).
(c) Time-stepping strategies (NS-SL).
Figure 3: Ablation of the proposed training strategies. a) The effect of edge masking on resolution invariance. EMP stands for edge masking probability. The y-axis shows the relative test error (at $t=t_{14}$) when the spatial resolution of the inputs and outputs varies from the training resolution ($92^{2}$). b) Autoregressive test errors ($\tau=\Delta t$) with and without fractional pairing fine-tuning. All values after the vertical line are considered as extrapolation in time. c) Autoregressive test errors ($\tau=2\Delta t$) with different time-stepping methods. All models are trained without the fractional pairing strategy and with $\tau_{\rm max}=8\Delta t$. All values after the vertical lines are considered as extrapolation in time.
On Temporal Processing with RIGNO.
RIGNO is designed to learn the solution operators of time-dependent PDEs and contains many elements in its design to enforce time continuity and enable accurate temporal approximation. To test these aspects, we focus on the NS-PwC and the NS-SL datasets (see SM Section D). The underlying trajectories are available from the PDEgym collection [12] on a time interval of $[0,1]$, sampled at equidistant time steps $t_{n}=n\Delta t$, with $\Delta t=0.05$ and $0\leq n\leq 20$. We trained RIGNO with data sampled at timesteps $\{t_{0},t_{2},t_{4},t_{6},t_{8},t_{10},t_{12},t_{14}\}$. Thus, we set $\tau=2\Delta t$ in our temporal marching procedure. During inference, we unroll RIGNO autoregressively with a time step size of $\Delta t$ and compare the test errors at all possible time indices $1\leq n\leq 20$. The resulting test errors are shown in Figure 3 (b,c). We have several observations to make about the results. i) The test errors grow in time as expected from an autoregressive rollout strategy. However, the growth of errors is not exponential but rather linear. ii) RIGNO is able to extrapolate well outside the temporal training domain. We see this from comparing the errors at time indices $n\geq 15$. These times have not been seen during training. Yet, RIGNO is able to generalize very well even to these unseen time points. iii) The proposed fractional pairing fine-tuning approach significantly enhances the accuracy with small unseen time steps. The errors are reduced by a significant factor, and in particular, fractional pairing helps generalize much better in the time extrapolation regime, $n\geq 15$, with significantly lower errors and lower growth rate of errors as compared to when it is switched off.
We also consider time marching strategies arising from the choice of parameters $\alpha,\beta$ in (11). We present the results with all the three choices of time stepping, i.e., output, residual and derivative stepping for the NS-SL dataset in Figure 3 (c). In addition to the slow error growth and good extrapolation beyond the times seen during training, we see that the derivative time-stepping strategy with $\alpha=1,\beta=\tau$ is clearly superior in this example to the other two, leading us to prefer it during our evaluations.
Resolution Invariance and RIGNO.
As mentioned before, a key requirement for a robust neural operator is to enforce an appropriate form of resolution invariance [2]. To test the abilities of RIGNO in this regard, we consider the NS-PwC dataset on random point clouds (see SM Section D). The training resolution for RIGNO is $92^{2}$. At inference, we test it on this and several other resolutions, presenting the test errors in Figure 3(a). We see from this figure that the resolution invariance errors with the base version of RIGNO are fairly small and indicate that the proposed local aggregation scheme enables RIGNO to generalize well across spatial resolution. Additionally, we observe that edge masking allows us to enforce resolution invariance to a greater extent. Moreover, the computational cost is roughly cut to half with edge masking probability $0.5$. As shown in Figure 2(b) and SM Figure G.1, using a larger model can compensate for the accuracy drop caused by edge masking.
Quantifying Model Uncertainties.
Not only does edge masking enable RIGNO to be (approximately) resolution invariant, it provides a simple method to quantify model uncertainties. As edge masking is random, invoking it during inference enables the model to provide multiple outputs for the same input. The mean of the outputs can be used as the model prediction as shown in SM Figure G.3. The same figure shows how the standard deviation of this ensemble of outputs is highly correlated with the test error. Thus, it can be used to provide an adaptive time marching strategy to reduce errors at inference, see SM Section G.4 for details. In SM Section G.5, we show that edge masking makes RIGNO more robust to noisy inputs as well.
4 Discussion.
In this paper, we present RIGNO, an end-to-end GNN-based neural operator for learning the solution operators of PDEs on arbitrary domains. Our model follows an encode-process-decode paradigm [38]. Given a point cloud on which the input data is defined (the physical mesh), the encoder is a GNN that maps the input data into a latent space defined on a regional mesh, which can be interpreted as a downsampled version of the underlying point cloud. The input point cloud can be very general, as no relational structure or graph must be specified on it. The processor maps the latent variables on the regional mesh into learned representations through multi-scale processing, implemented through multiple stacked message-passing neural networks. Finally, the decoder transforms the processed variables from the regional mesh to physical variables on the output point cloud. Again, the only graph in this layer connects the regional mesh nodes to the points in the output cloud. No relational structure or graph needs to be defined on the output point cloud, making RIGNO very general in terms of its abilities to handle data on arbitrary point clouds.
In addition to this basic construction (see Figure 1), which was inspired by [16], RIGNO introduces many novel elements to satisfy two important requirements. First, as desired in [15, 2], (approximate) resolution invariance is essential for structure-preserving operator learning. RIGNO uses overlaps between the sub-regions in the regional mesh connecting points in the physical mesh and more importantly, edge masking, to enable approximate resolution invariance, a feature which is hard to achieve for neural operators on arbitrary domains. Second, RIGNO is tailored to process time-dependent PDEs. In this context, temporal continuity and control on error growth over time are key properties that the neural operator needs to uphold. Here, we have proposed many elements in RIGNO to achieve these requirements, namely lead time conditioning, all2all training, and a fractional pairing fine-tuning strategy. This allows us to maximize training steps for a given number of trajectories, while allowing for general rollout strategies at inference.
We have tested RIGNO on a suite of challenging problems including PDEs with data on structured/unstructured meshes, random point clouds, and Cartesian grids. RIGNO is shown to be significantly more accurate than baselines designed for such arbitrary domains. We also observe that RIGNO is more accurate than state-of-the-art neural operators such as CNO, FNO and scOT, even for PDEs on Cartesian grids. The gains in accuracy are not marginal, and can be of an order of magnitude on many problems, demonstrating the expressive power of the proposed architecture. We show that RIGNO scales with both model and data size. Simultaneously, RIGNO has excellent resolution invariance and temporal continuity properties. It also generalizes very well to interpolated and extrapolated time levels, validating novel elements such as fractional pairing that were introduced into the model to improve temporal processing.
Importantly, we must also highlight some limitations of RIGNO and indicate avenues for further extension. Here, we have focused on PDEs in two space dimensions. In principle, RIGNO can readily be extended to PDEs in three space dimensions, although the associated computational cost and paucity of publicly available 3D benchmarks currently limits our exploration of this crucial direction. Similarly, our focus here was on presenting the model and extensive empirical investigation. Important theoretical questions such as those of universal approximation of PDE solution operators with RIGNO will be considered in future work. Similarly, we have empirically demonstrated that RIGNO scales with the number of training samples; however, proving a rigorous generalization error bound, such as that of [17], merits further analysis. Moreover, the scaling curves of Figure 2 suggest that a large number of training samples are necessary for neural operators to achieve small error. This issue is common to all standalone neural operators [12] and can be alleviated by building foundation models such as those proposed in [12, 11, 41, 29]. Using RIGNO as the backbone for a foundation model is another avenue for further exploration. Finally, downstream applications of RIGNO to tasks such as uncertainty quantification (UQ), inverse problems and PDE-constrained optimization will be considered in future work.
Acknowledgements
The contribution of S. Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191).
Supplementary Material for RIGNO: A Graph-based Framework
for robust and accurate operator learning for PDEs on arbitrary domains.
Appendix A Graph Construction
A.1 Encoder and decoder edges
The edges between the physical and regional nodes in the encoder and in the decoder are defined based on support neighborhoods around each regional node. We consider a circular neighborhood around each regional node (see Figures 1 and 1(a)) and calculate the radii of the circles such that the union of all the circles covers the whole domain. To achieve this, we first construct a Delaunay triangulation on the regional nodes. We then calculate all medians of all the simplices in the triangulation. The minimum radius of a node is then calculated as two-thirds of the longest median that crosses that node. The minimum radius is then multiplied by an overlap factor, which is a hyperparameter of the model. A similar method can be applied to three-dimensional point clouds as well. Each regional node is then connected to all the physical nodes that lie inside its corresponding support sub-region. With a larger overlap factor, the regional nodes receive information from more physical nodes in the encoder. Too large overlap factors might be unnecessary since regional nodes are assumed to encode information locally. However, in the decoder, overlap factor plays a more important role. With minimum overlaps, many physical nodes will end up being connected to only a single regional node, which might result in overly biased decoded latent features. Therefore, we generally use different overlap factors in the encoder and in the decoder.
A.2 Regional nodes
By defining the edge connections between the regional and the physical nodes based on support sub-regions in the encoder and in the decoder, RIGNO allows for unstructured physical and regional meshes. Within this framework, even for an unstructured set of physical nodes, the regional nodes can still follow a structure (e.g., lie on a uniform grid) since the choice of the regional nodes is completely arbitrary. However, constructing an unstructured regional mesh is favorable since local resolutions of the physical mesh can be inherited by the regional mesh. To this end, we always define the regional nodes by randomly sub-sampling the given physical nodes by a factor that is given as a hyperparameter. The set of regional nodes may be different during inference or in each training epoch.
The sub-sampling factor is a fixed hyperparameter, which makes the resolution of the regional mesh dependent on the resolution of the physical mesh. If a model is only trained with a fixed mesh resolution, changing the resolution at inference might affect its accuracy. However, since the choice of the regional mesh nodes is completely arbitrary, the resolution of the regional mesh can be corrected in order to correspond to the resolution of the training regional meshes. For zero-shot super-resolution inference, this can be easily done by sub-sampling the physical mesh with a higher sub-sampling factor. For zero-shot sub-resolution inference, however, a suitable mesh-refinement algorithm must be employed. In this work, we refine the mesh by first constructing a standard triangulation, randomly picking the necessary number of triangles, and adding the centroid of those triangles to the regional mesh. With this method, we ensure that local resolutions are inherited from the original mesh.
A.3 Processor edges
The processor of RIGNO consists of a sequence of message-passing blocks with skip connections, which act on the regional nodes and the edges that connect them. We define the first level of the bidirectional edges between the regional nodes as all the unique pairs of connected vertices in a Delaunay triangulation on the regional nodes. In addition to the first level edges, we also define edges that connect long-distance regional nodes, which allows learning multiple spatial scales and enables a more efficient propagation of information in a message-passing step. The construction of these edges is illustrated in Figure 1 as multi-scale edges. For the second level, we randomly choose a subset the regional nodes and repeat the above process on them. We repeat this process multiple times by sub-sampling the subsets again.
The structure of the regional mesh can be regarded as a hierarchy of multiple sub-meshes with different resolutions and shared nodes (see Figures 1(b) and 1(c), and the multi-scale edges box of Figure 1). In one message-passing step, the coarser sub-meshes, which have longer edges, allow the information to flow in larger scales, while the finer sub-meshes allow for assimilating small-scale details. All regional nodes are part of the finest sub-mesh, while only a few nodes take part in the coarsest one. In a single message-passing step, it is not possible for all the nodes to receive information from longer distances. Because of this, it is essential to have multiple message-passing steps in the processor to allow for the information to propagate through the whole domain.
A.4 Fixed and Variable physical meshes
For some applications in engineering (e.g., shape optimization), the physical mesh changes for every sample in the training dataset or at inference. Since the coordinates of the input discretization will change, the graph construction needs to be carried out every time. For training with large datasets, these graphs can be constructed only once as a pre-processing step. At inference, however, the time for constructing the graph is considerable compared to the execution of the model. On the other hand, another class of applications in science and engineering only require training and inference on a fix unstructured physical mesh. For these problems, constructing the graph can only be done once.
For the latter class of problems, RIGNO is trained using a fixed physical mesh. However, if the graph construction is only done once during training, regional mesh will consequently be fixed too. In this case, although the encoder and the decoder are invariant to the input discretization, the processor might be biased to that specific fixed regional mesh. Such a model can exhibit a poor performance when it is inferred with a different regional mesh. In order to avoid this situation, we update the regional mesh before every epoch during training.
(a) Support sub-regions of regional nodes.
(b) Level 1 regional mesh edges.
(c) Level 2 regional mesh edges.
(d) Wrapped regional mesh edges.
Figure A.1: Construction of regional mesh edges and support sub-regions with periodic boundary conditions. The dashed lines show the domain boundaries. Large and small purple dots represent the main regional nodes and the ghost regional nodes, respectively. Black lines represent the main edges, and the purple lines the cross-boundary edges. Note that in Figure 1(a), the parts of the domain that are close to the domain boundaries and do not lie in any circle are covered by at least one regional node on the other side of the boundary.
Appendix B Boundary Conditions in RIGNO
B.1 Periodic boundary conditions
The majority of the problems considered in this work have periodic boundary conditions. The flexibility of graphs allows us to enforce periodic boundary conditions in the architecture without any significant computational overhead. To achieve this, we make the following three changes to our framework for problems with periodic boundary conditions: 1) Initialize the node features with a periodic encoding of the coordinates instead of raw coordinates; 2) extend the domain before defining the encoder and the decoder edges; and 3) extend the domain before defining the processor edges.
In order to motivate the above changes, let us consider two nodes, one close to the left and one close to the right boundary. With (normalized) raw coordinates, the initial features of these two nodes will be very different as one will take a value close to 00 and the other one takes a value close to $1$. With periodic boundary conditions, however, these two points will be close to each other if the frame of the reference domain is slightly shifted. The invariance to a shift in the frame of reference means that these two nodes should be connected in the graphs. In simpler words, we assume during graph construction that the both ends of a domain with periodic boundary conditions are connected. In one-dimensional settings, this set-up corresponds to considering a circular domain, see Figure B.1 for an illustration. For a square two-dimensional domain, connecting the two ends in both direction corresponds to a torus.
Inspired by the ghost node technique in classical numerical methods, we extend the domain in all directions by constructing a 3$\times$3 tiled structure, where each tile contains the repeated coordinates of the given point cloud. All the Delaunay triangulations in the graph construction phase are built based on the extended domain. Figure A.1 shows the resulting support sub-regions, as well as the extra edges that are introduced because of the domain extensions for a minimal set of regional nodes. We refer to these extra edges as cross-boundary edges. Notice how all the white areas in Figure 1(a) are covered by at least on circle on the other side of the domain. In order to avoid introducing extra costs to the architecture, the ghost nodes are discarded after graph construction and only the cross-boundary edges are kept. In order to keep the extra edges, we replace each ghost node with its corresponding node in the original point cloud. The result is illustrated in Figure 1(d).
Figure B.1: Graph structure for periodic boundary conditions in one-dimensional settings. The green dots represent physical nodes, the purple circles represent regional nodes, and the solid lines represent edge connections. Note that only a few physical nodes are visualized. The bottom right part of the figure partly shows the receptive fields of two regional nodes in the encoder. Here, the red lines show the edges that are randomly masked, and hence will not be used in that particular message passing block.
B.2 General boundary conditions
In problems with general (non-periodic) boundary conditions, we rely on the capacity of the model to learn the boundary conditions implicitly from data. Although this seems to be sufficient for the problems considered in this work, we would like to highlight that more sophisticated mechanisms are needed for non-homogeneous boundary conditions or homogeneous boundary conditions that change in each sample. One possible approach is to consider a separate type of edge or node for encoding data on the boundaries, similar to the type extension method that is proposed in [32], to differentiate between mesh edges and proximity-based edges for Lagrangian systems. In this work, however, we only consider one node type and one edge type, since the considered solution operators are not dominated by boundary conditions.
Appendix C Temporal Processing
C.1 Time Stepping Strategies
As described in the main text, we consider three time stepping strategies, namely direct output stepping, residual stepping, and time-derivative stepping. It is important to normalize the outputs correspondingly in each of these strategies since the nature and the statistics of the target variables are different in each case. Therefore, we pre-compute the appropriate statistics in each strategy based on the training samples and use them to normalize and de-normalize the inputs and the outputs. With direct output stepping, the statistics of the inputs and the outputs are trivial. For residual stepping, we pre-compute the statistics of the residuals with all the lead times that are present in the training. For time-derivative stepping, we similarly consider all lead times and compute the statistics of the residual divided by the lead time. Figure C.1 illustrates flowcharts for input and output normalization with each stepping strategy.
(a) Output stepping.
(b) Residual stepping.
(c) Time-derivative stepping.
Figure C.1: Flowcharts of input and output normalization strategies. Blue boxes represent raw data and green boxes represent normalized data. We apply a min-max normalization for the input time $t$, which is omitted in the flowcharts for simplicity. The loss is computed on normalized values while the errors are computed on the raw targets and de-normalized estimates. Note that the raw target and the de-normalized estimates may be normalized in a different manner or with different statistics (e.g., global equation-wide statistics as in [12]) before the error is computed on them.
C.2 Temporal Inference
Given any set of partial lead times $\{\tau_{i}\}_{i=1}^{n}$ that sum up to a time $t\in\Omega_{t}$, a time-continuous operator $\mathcal{G}_{\theta}$ can be applied sequentially to reach the solution at time $t$:
$u(t)\simeq\mathcal{G}_{\theta}\left(\mathcal{G}_{\theta}\left(\mathcal{G}_{\theta}\left(\dots,c,\sum_{i=1}^{n-3}\tau_{i},\tau_{n-2}\right),c,\sum_{i=1}^{n-2}\tau_{i},\tau_{n-1}\right),c,\sum_{i=1}^{n-1}\tau_{i},\tau_{n}\right).$ (13)
This gives an infinite number of ways to infer the learned operator for reaching $u(t)$. However, since $\mathcal{G}_{\theta}$ is not exact, the approximation errors generally accumulate and can lead to very high errors for long-time estimations. With larger lead times it is possible to decrease the number of times that the operator is applied or to completely avoid autoregressive inference by directly using $\tau=t$.
In all the time-dependent datasets considered in this work, for simplicity of exposition, we have considered $N_{t}=20$ or $N_{t}=21$ time snapshots $\{t_{n}\}_{n=1}^{N_{t}}$. In all trainings, only snapshots with even indices up to $t_{14}$ are used. Snapshots with $t>t_{14}$ are reserved for time extrapolation tests, and snapshots with $t<t_{14}$ with even indices are reserved for time interpolation tests. Error at snapshot $t_{14}$ is used as a reference for the benchmarks. Since the trained operator is continuous in time and can be inferred with different lead times, there are infinitely many ways to reach the solution at $t_{14}$ via autoregressive inference of the trained operator. The optimal autoregressive path can be different for different operators. Figure C.2 illustrates the training and test time snapshots as well as different autoregressive paths for reaching $t_{14}$. In the presented benchmarks in Table 1, we calculate the error with 3 different autoregressive paths: i) $\{\tau_{i}=2\Delta t\}_{i=1}^{7}$ which is denoted by AR-2; ii) $\{\tau_{1}=4\Delta t,\tau_{2}=4\Delta t,\tau_{3}=4\Delta t,\tau_{4}=2\Delta t\}$ which is denoted by AR-4; and iii) $\{\tau_{1}=14\Delta t\}$, denoted by DR. We always report the lowest error among these three rollout paths.
Figure C.2: Snapshots used for training, time interpolation, and time extrapolation. The in-distribution and extrapolation evaluation snapshots are marked with thick borders. The lower arrows show the autoregressive steps for evaluating a learned operator at $t_{14}$ with $\tau_{\rm max}=\Delta t$ (14 steps). Note that in this rollout path, a lead time ($\Delta t$) that is smaller than the time resolution of the training dataset ($2\Delta t$) is applied 14 times in the evaluation. The upper arrows show the autoregressive steps for evaluating a learned operator at $t_{20}$ with $\tau_{\rm max}=4\Delta t$ (5 steps).
C.3 Fractional Pairing
We use the fractional pairing strategy to achieve a truly time-continuous operator that can be inferred with lead times smaller than the smallest time difference in the training dataset. In fractional pairing, we approximate the solution at an intermediary time by inferring a model that has been trained on the original time resolution. This approximation is then given as input to the same model. By choosing a suitable intermediary time, we can have an input-output pair that is not present in the training dataset and have a smaller lead time than the time resolution of the training dataset. This idea is illustrated in Figure C.3. With this, we can either calculate the loss gradients for both of the inferences, or we can cut the gradient at the intermediary time and treat the intermediary solution as a normal model input, which is referred to as GradCut (see bottom left of Figure C.3). In the experiments of this work, we always use fractional pairing with GradCut.
If the approximation at the intermediary time is not accurate enough, fractional pairing strategy can lead to training instabilities. Therefore, it is recommended to use this technique only as a fine-tuning stage. Furthermore, since the approximations are made with fractional lead times, i.e., lead times that are not present in the training dataset, it is also essential to employ other mechanisms that help generalize to fractional lead times with regular training strategies, such as consistent time stepping strategies and time-conditioned normalization layers. A further consideration that helps the stability of this strategy is avoiding smaller-than-time-resolution lead times in the first forward pass. Such small lead times can be seen as an extrapolation in the seen distribution of lead times during training. It has been observed that even without fractional pairing, the models generalize better to interpolated lead times than they do to extrapolated lead times. Therefore, we can approximate the intermediary solution with an interpolated lead time, and use a smaller lead time only in the second forward pass. In Figure C.3, this corresponds to only using the two right-most additional pairs in the green box. Note that none of the additional pairs of the orange box will be viable in this setting.
Figure C.3: Schematic of the fractional pairing strategy. The additional pairs obtained by fractional pairing correspond to smaller lead times than the smallest time difference in the original trajectory. On the bottom left of the figure, the forward passes (solid black arrows) and the backward gradient flows (dashed arrows) are illustrated for an existing pair $(t_{m},t_{n})$. With the GradCut setting, the gradient flow is stopped at $t_{\ast}$.
Appendix D Datasets
RIGNO is tested on a suite of 12 datasets [31] which covers operators arising from compressible Euler (CE) equations of gas dynamics, incompressible Navier-Stokes (NS) equations of fluid dynamics, wave equation (WE), heat equation (HE) of thermal dynamics, Allen-Cahn equation (ACE), and force balance equation in solid mechanics (FB). Table D.1 summarizes the main characteristics of all the datasets. A random sample of each dataset is visualized in Figures D.2-D.13. This collection of datasets covers diverse multi-scale, nonlinear, complex phenomena in systems arising from PDEs such as shock waves, shear layers, mixing, and turbulence. This makes the effort to learn the underlying solution operator from data worthwhile for these datasets; since running simulations of these systems with any classical numerical method would require extensive computational resources. Moreover, by considering these datasets, we cover a diverse set of domains and meshes from Cartesian uniform grids on a unit square to structured and unstructured meshes on fixed and varying irregular geometries.
The problems in the 8 datasets from [12] are all solved in the unit spatio-temporal domain $\Omega_{t}\times\Omega_{x}=[0,T]\times[0,1]^{2}$. The final time is $T=1$ except for the ACE dataset where $T=0.0002$. For each dataset, a parameterized distribution of initial conditions is considered, and initial conditions are randomly drawn by sampling those distributions. The distribution of the initial conditions characterizes the dataset, the level of complexity of the dynamics captured in it, the phenomenological features exhibited in the solution function, and the corresponding solution operator. In all these datasets, $N_{t}=21$ snapshots of each sample are collected at times $\Omega_{t}^{\Delta}=\{t_{n}=n\Delta t\}_{n=0}^{N_{t}-1}$ with uniform spacing $\Delta t$; except for the ACE and the Wave-Layer datasets where $N_{t}=20$. Each snapshot originally has a 128$\times$128 resolution, with uniform discretization in both spatial directions. In order to test the baselines on a random point cloud with these datasets, we randomly shuffle the coordinates and only keep the first 9216 coordinates in our experiments. Figure D.1 illustrates a few random point clouds obtained from this procedure.
Figure D.1: Random point clouds (9216 points) obtained from a 128$\times$128 uniform grid.
Table D.1: Abbreviations and main characteristics of all datasets. IC: initial condition; BC: boundary conditions.
Abbreviation PDE BC Domain Mesh Source Visualization
Heat-L-Sines HE Dirichlet L-shape unstructured generated Figure D.2
Wave-C-Sines WE Dirichlet circle unstructured generated Figure D.3
NS-Gauss INS periodic square uniform grid [12] Figure D.4
NS-PwC INS periodic square uniform grid [12] Figure D.5
NS-SL INS periodic square uniform grid [12] Figure D.6
NS-SVS INS periodic square uniform grid [12] Figure D.7
CE-Gauss CE periodic square uniform grid [12] Figure D.8
CE-RP CE periodic square uniform grid [12] Figure D.9
ACE ACE periodic square uniform grid [12] Figure D.10
Wave-Layer WE absorbing square uniform grid [12] Figure D.11
AF CE various airfoil structured [21] Figure D.12
Elasticity FB various holed square unstructured [21] Figure D.13
D.1 Equations
D.1.1 Incompressible Navier-Stokes equations (INS)
The incompressible Navier-Stokes equations are given by
[…]
where $\boldsymbol{v}:\Omega_{t}\times\Omega_{x}\to\mathbb{R}^{2}$ is the velocity field with components $(v_{x},v_{y})^{\top}$, $p:\Omega_{t}\times\Omega_{x}\to\mathbb{R}_{+}$ is the pressure, and $\nu\in\mathbb{R}$ is the fluid viscosity. In all the INS datasets, periodic boundary conditions are considered and a small viscosity $\nu=4\times 10^{-4}$ is applied to the high Fourier modes only.
D.1.2 Compressible Euler equations (CE)
The compressible Euler equations in the conservational form are given by
$\partial_{t}\boldsymbol{u}+\nabla\cdot\boldsymbol{F}$ $=\boldsymbol{S},$ (15)
$\boldsymbol{u}$ $=(\rho,\rho\boldsymbol{v},E)^{\top},$
$\boldsymbol{F}$ $=(\rho\boldsymbol{v},\rho\boldsymbol{v}\otimes\boldsymbol{v}+p\boldsymbol{I},(E+p)\boldsymbol{v})^{\top},$
$E$ $=\frac{1}{2}\rho\|\boldsymbol{v}\|^{2}+\frac{p}{\gamma-1},$
$\boldsymbol{S}$ $=(0,-\rho,0,-\rho v_{x})\frac{\partial\phi}{\partial x}+(0,0,-\rho,-\rho v_{y})\frac{\partial\phi}{\partial y},$
with $\gamma=1.4$, where $\boldsymbol{F}$ is the flux matrix, $E$ is the total energy density (energy per unit volume), and $\phi$ is the gravitational potential field. The gravity source term is not considered in any of the datasets of this work.
D.1.3 Allen-Cahn equation (ACE)
The Allen–Cahn equation describes phase transition in materials science and is given by
$\partial_{t}u-\nabla^{2}u={\epsilon^{2}}(u^{3}-u),$ (16)
where $\epsilon$ is the reaction coefficient.
D.1.4 Wave equation (WE)
The wave equation is considered with a fixed-in-time propagation speed $c$ that depends on the spatial location. This setting describes the propagation of acoustic waves in an inhomogeneous medium:
$\partial_{tt}u-c^{2}\nabla^{2}u=0.$ (17)
The wave equation can be written in the form of (1) by considering the first time derivative $\partial_{t}u$ as a variable and writing it as a system of equations. The propagation speed can either be added as an input to the differential operator, or as another variable with $\partial_{t}c=0$. In the former case, it should also be added to the inputs of any solution operator.
D.1.5 Heat equation (HE)
The heat equation describes the diffusion of heat in a medium and is given by
$\partial_{t}u\;=\;a^{2}\nabla^{2}u,$ (18)
where $a$ is the diffusivity constant.
D.1.6 Solid body force balance equation (FB)
The force balance equation of a solid body is given by
$\rho\partial_{t}\boldsymbol{u}+\nabla\cdot\boldsymbol{\sigma}=0,$ (19)
where $\rho$ denotes mass density, $\boldsymbol{u}$ the displacement field, and $\boldsymbol{\sigma}$ the stress tensor. The strain field is related to the displacement field via kinematic relations. The system is closed by considering a constitutive model which describes the relation between strain and stress.
D.2 Dataset details
D.2.1 Heat-L-Sines
This dataset considers time-dependent solutions of the heat equation over an L-shaped domain which is formed by removing the rectangular portion [1, 1.5]×[0.5, 1]11.50.51[1,\,1.5$\times$[0.5,\,1][ 1 , 1.5 ] × [ 0.5 , 1 ] from the larger rectangle [0.5, 1.5]×[0, 1]0.51.501[0.5,\,1.5$\times$[0,\,1][ 0.5 , 1.5 ] × [ 0 , 1 ]. We consider constant diffusivity $a=1.0$ and zero homogeneous Dirichlet boundary conditions. We prescribe a superposition of sinusoidal modes as the initial condition:
$u_{0}(x_{1},x_{2},\boldsymbol{\mu})\;=\;-\tfrac{1}{M}\sum_{m=1}^{M}\mu_{m}\,\frac{\sin\bigl{(}\pi mx_{1}\bigr{)}\,\sin\bigl{(}\pi mx_{2}\bigr{)}}{m^{0.5}},$ (20)
where $\boldsymbol{\mu}\in[-1,1]^{d}$, $M=10$, and each $\mu_{m}$ is sampled independently and uniformly from $[-1,1]$. A Delaunay-based triangular mesh with characteristic length $0.008$ is used, yielding about $14{,}047$ nodes and $27{,}590$ elements. We simulate the heat equation for $100$ time steps with a time step size of $\Delta t=5\times 10^{-5}$ and keep $21$ (including the initial condition) equally spaced snapshots from each trajectory.
Figure D.2: Snapshots of solution trajectories of the Heat-L-Sines dataset.
D.2.2 Wave-C-Sines
This dataset pertains to time-dependent solutions of the two-dimensional wave equation over a circular domain of radius $1$, centered at (0.5, 0.5)0.50.5(0.5,\,0.5)( 0.5 , 0.5 ). We set $c=2.0$, constant in time and in space, and impose homogeneous Dirichlet boundary conditions ($u=0$) on the boundary of the circle. The initial condition for $u$ at $t=0$ is specified as a superposition of sinusoidal modes:
$u_{0}(x,y)\;=\;\frac{\pi}{M^{2}}\sum_{i,j}^{M}a_{ij}\,(i^{2}+j^{2})^{-r}\,\sin(\pi ix)\,\sin(\pi jy),$ (21)
where $K=10$, $r=0.5$, and $a_{ij}\sim\mathcal{U}_{[-1,1]}$. We employ a Delaunay-based triangular mesh with characteristic length $0.015$, resulting in roughly $16{,}431$ nodes and $32{,}441$ elements. A time step of $\Delta t=0.001$ is used, and solutions are computed for $100$ time steps via a finite element solver. We then downsample the computed solution fields in time, keeping $21$ (including the initial condition) equally spaced time frames, which constitute the dataset.
Figure D.3: Snapshots of solution trajectories of the Wave-C-Sines dataset.
D.2.3 NS-Gauss
The initial velocities are obtained from an initialized vorticity defined by $\omega(x,y)=\partial_{x}v_{y}-\partial_{y}v_{x}$ by using the incompressibility condition [28]. The vorticity field is initialized as a linear superposition of $M=100$ Gaussian distributions:
$\omega^{0}(x,y)=\sum_{i=1}^{M}\frac{\alpha_{i}}{\sigma_{i}}\exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (22)
with $\alpha_{i}\sim\mathcal{U}_{[-1,1]}$, $\sigma_{i}\sim\mathcal{U}_{[0.01,0.1]}$, $x_{i}\sim\mathcal{U}_{[0,1]}$, and $y_{i}\sim\mathcal{U}_{[0,1]}$. A random sample of this dataset is visualized in Figure D.4.
Figure D.4: Snapshots of solution trajectories of the NS-Gauss dataset. Each row corresponds to one variable.
D.2.4 NS-PwC
The initial velocities are obtained from an initialized vorticity by using the incompressibility condition. The initial vorticity $\omega^{0}$ of the flow field is initialized as a collection of piece-wise constant squares
$\omega^{0}(x,y)=c_{i,j}\mbox{ in }[x_{i-1},x_{i}]\times[y_{j-1},y_{j}],$ (23)
with $x_{i}=y_{i}=\frac{i}{P}$ for $i=0,1,2,\dots,P$, and $c_{i,j}\sim\mathcal{U}_{[-1,1]}$. The number of squares in each direction is $P=10$, resulting in $100$ squares in total. A random sample of this dataset is visualized in Figure D.5.
Figure D.5: Snapshots of solution trajectories of the NS-PwC dataset. Each row corresponds to one variable.
D.2.5 NS-SL
The velocity field $\boldsymbol{v}=(v_{x},v_{y})$ is initialized as
$v^{0}_{x}(x,y)$ =$=\begin{cases}\tanh\left(2\pi\frac{y-0.25}{\rho}\right)&\mbox{ for }y+\sigma_{\delta}(x)\leq\frac{1}{2}\\ \tanh\left(2\pi\frac{0.75-y}{\rho}\right)&\mbox{ otherwise}\end{cases},$ (24)
$v^{0}_{y}(x,y)$ $=0,$
where $\sigma_{\delta}:[0,1]\to\mathbb{R}$ is a perturbation of the initial data given by
$\sigma_{\delta}(x)=\xi+\delta\sum_{k=1}^{p}\alpha_{k}\sin(2\pi kx-\beta_{k}).$ (25)
The parameters are sampled as $p\sim\mathcal{U}_{\{7,8,9,10,11,12\}}$, $\alpha_{k}\sim\mathcal{U}_{[0,1]}$, $\beta_{k}\sim\mathcal{U}_{[0,2\pi]}$, $\delta=0.025$, $\rho\sim\mathcal{U}_{[0.08,0.12]}$, and $\xi\sim\mathcal{U}_{[-0.0625,0.0625]}$. A random sample of this dataset is visualized in Figure D.6.
Figure D.6: Snapshots of solution trajectories of the NS-SL dataset. Each row corresponds to one variable.
D.2.6 NS-SVS
The vorticity of the flow is initialized as
$\omega^{0}(x,y)=\phi_{\rho}(x,y)*\omega^{{}^{\prime}}(x,y),$ (26)
where
$\omega^{{}^{\prime}}(x,y)$ $=\delta(x-\Gamma)-\int_{\mathbb{T}^{2}}\!\,\mathrm{d}\Gamma,$ (27)
$\phi_{\rho}(x,y)$ $=\rho^{-2}\psi\left(\frac{\|x\|}{\rho}\right),$ (28)
$\psi(r)$ $=\frac{80}{7\pi}\left[(r+1)_{+}^{3}-4(r+\frac{1}{2})_{+}^{3}+6r_{+}^{3}-4(r-\frac{1}{2})_{+}^{3}+(r-1)_{+}^{3}\right],$ (29)
$\Gamma$ =$=\{(x,y)\in\mathbb{T}^{2}\mid y=\frac{1}{2}+0.2\sin(2\pi x)+\sum_{i=1}^{10}\alpha_{i}\sin(2\pi(x+\beta_{i}))\}.$ (30)
The random variables $\alpha_{i}$ and $\beta_{i}$ are given by $\alpha_{i}\sim\mathcal{U}_{[0,0.003125]}$, $\beta_{i}\sim\mathcal{U}_{[0,1]}$. The parameter $\rho$ is fixed as $\rho=\frac{5}{128}$. A random sample of this dataset is visualized in Figure D.7.
Figure D.7: Snapshots of solution trajectories of the NS-SVS dataset. Each row corresponds to one variable.
D.2.7 CE-Gauss
The $\omega$ field is initialized as a linear combination of $M=100$ Gaussians
$\omega^{0}(x,y)=\sum_{i=1}^{M}\frac{\alpha_{i}}{\sigma_{i}}\exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (31)
where $\alpha_{i}\sim\mathcal{U}_{[-1,1]}$, $\sigma_{i}\sim\mathcal{U}_{[0.01,0.1]}$, $x_{i}\sim\mathcal{U}_{[0,1]}$, and $y_{i}\sim\mathcal{U}_{[0,1]}$. The density and pressure are initialized with $\rho=0.1$ and $p=2.5$, both constant in the whole domain. Similar to the NS-Gauss dataset, the initial velocity field is obtained from the initialized vorticity by using the incompressibility condition. A random sample of this dataset is visualized in Figure D.8.
Figure D.8: Snapshots of solution trajectories of the CE-Gauss dataset. Each row corresponds to one variable.
D.2.8 CE-RP
This dataset corresponds to the four-quadrant Riemann problem, which generalizes the standard Sod shock tube to two spatial dimensions. The domain is divided into a grid of $P\times P$ square subdomains $D_{i,j}=\{(x,y)\in\mathbb{T}^{2}\mid\frac{i-1}{P}\leq x<\frac{i}{P},\frac{j-1}{P}\leq y<\frac{j}{P}\}$. The solution on each of these subdomains is constant and initialized as
$(\rho,v_{x},v_{y},p)(x,y)=(\rho_{i,j},u_{i,j},v_{i,j},p_{i,j}),$ (32)
where $\rho_{i,j}\sim\mathcal{U}_{[0.1,1]}$, $u_{i,j}\sim\mathcal{U}_{[-1,1]}$, $v_{i,j}\sim\mathcal{U}_{[-1,1]}$, and $p_{i,j}\sim\mathcal{U}_{[0.1,1]}$. In this dataset, $P=2$ is considered, which constitutes 4 subdomains in total. A random sample of this dataset is visualized in Figure D.9.
Figure D.9: Snapshots of solution trajectories of the CE-RP dataset. Each row corresponds to one variable.
D.2.9 ACE
The Allen–Cahn equation (16) are considered with periodic boundary conditions and reaction coefficient $\epsilon=220$. $u$ is initialized as
$u^{0}(x,y)=\frac{1}{M^{2}}\sum_{i,j=1}^{M}a_{ij}\cdot(i^{2}+j^{2})^{-r}\sin(\pi ix)\sin(\pi jy),$ (33)
where $M$ is a random integer randomly sampled from $[16,32]$, $r\sim\mathcal{U}_{[0.7,1]}$, and $a_{ij}\sim\mathcal{U}_{[-1,1]}$. A random sample of this dataset is visualized in Figure D.10.
Figure D.10: Snapshots of solution trajectories of the ACE dataset. Each row corresponds to one variable.
D.2.10 Wave-Layer
The wave equation (17) is considered with the fixed-in-time propagation speed $c(x,y)$ as a piece-wise constant function in a vertically layered medium with $n\in\{3,4,5,6\}$ layers and randomly defined intersections. More details about the definition of layers are given in SM B.2.11 of [12]. $u$ is initialized as a linear combination of $n\in{2,3,4,5,6}$ Gaussian distributions centered at random locations in the domain. A random sample of this dataset is visualized in Figure D.11.
Figure D.11: Snapshots of solution trajectories of the Wave-Layer dataset. Each row corresponds to one variable.
D.2.11 AF
In this dataset, the transonic flow is considered over different airfoils. The steady-state form of the compressible Euler equations (15) govern this system. The viscous effects are ignored and the far-field boundary conditions of $\rho_{\infty}=1$ , $p_{\infty}=1.0$, $M_{\infty}=0.8$ and zero angle of attack are considered. At the airfoil boundary, no-penetration conditions are considered. We refer the reader to [21] for more details about how the airfoil geometries are drawn. The flow is computed on a C-grid mesh with 200$\times$50 quadrilateral elements by a second-order implicit finite volume solver. The original dataset is enhanced by adding a distance field from the boundary of the airfoil which is used as the input of the model. The target value used in this work is the density $\rho$.
Figure D.12: Solution samples of the AF dataset (only density). The full domain is much larger and the mesh is more sparse than the visualized frame around the airfoil.
D.2.12 Elasticity
The solid body force balance equation (19) is considered for a unit square hyper-elastic incompressible Rivlin-Saunders specimen with a hole at its center. The geometry of the hole is randomly sampled from a certain distribution (see [21] for more details) such that the radius always takes a value between $0.2$ and $0.4$. The specimen is clamped at the bottom boundary (zero displacement) and is under a constant vertical tension traction on its top boundary. The stress field is computed for different geometries with a finite elements solver with about $100$ quadratic quadrilateral elements. The target function in this dataset is the stress field, which is available at $972$ unstructured coordinates. The original dataset is enhanced by adding a distance field from the boundary of the hole that is given as input to the models.
Figure D.13: Solution samples of the Elasticity dataset.
Appendix E Details of the Experiments
The hyperparameters of RIGNO-18 and RIGNO-12 are listed in Table E.1. The only difference between the two architectures is in the number of message passing blocks. The presented settings in this section are used in all the presented experiments, unless otherwise stated. A JAX [3] implementation of RIGNO is available on Github (https://github.com/camlab-ethz/rigno).
The last feed-forward block in the decoder that independently acts on physical nodes does not have normalization layers, i.e., it is only composed of a two-layer MLP. A nonlinear activation function is applied on the output of the hidden layer in all MLPs. For the main MLPs of the feed-forward blocks, we use the swish activation function [36]; whereas for the MLPs of the conditioned-normalization layers, the sigmoid function is used.
Table E.1: Architectural hyperparameters of RIGNO and their corresponding values for RIGNO-18 and RIGNO-12.
RIGNO-18 RIGNO-12 Description
2.7M 1.9M Total number of trainable parameters.
18 12 Number of message passing blocks.
4 4 Factor for sub-sampling the given coordinates. As an example, with 16,000 input coordinates, we get 4,000 coordinates in the regional mesh.
1.0 1.0 Overlap factor of the support sub-regions in the encoder.
2.0 2.0 Overlap factor of the support sub-regions in the decoder.
6 6 Number of regional edge levels.
128 128 Dimension of the latent node features.
128 128 Dimension of the latent edge features and messages.
1 1 Number of the hidden layers in the MLP of the feed-forward blocks.
128 128 Dimension of the hidden layers in the MLP of the feed-forward blocks.
16 16 Dimension of the hidden layers in the MLPs of conditional normalization layers.
4 4 Number of frequencies for encoding coordinates, i.e., $K_{\rm freq}$ in Equation (4).
0.5 0.5 Probability for edge masking.
We train our models with 1024 training samples for all datasets; except for the AF dataset where 2048 samples are used in the trainings. We use a different set of 128 samples for validation, and 256 samples for testing. The test errors reported in the second part of Table 1 (uniform grid datasets) are exceptionally calculated on 240 test samples. We train our models for 2500 epochs on the time-dependent datasets, and for 3000 epochs on the time-independent datasets. We calculate the errors on the validation dataset regularly and keep the model parameters that give the lowest validation error. We use batch size 16, the AdamW [26] optimizer with weight decay $10^{-8}$, the mean squared error loss function, and cosine learning rate decay with linear warm-up period. The learning rate is warmed up in the first 2% training epochs from $10^{-5}$ to $2\times 10^{-4}$, and decays with cosine schedule to $10^{-5}$ in the next 88% of the epochs. In the last 10% epochs, it exponentially decays to $10^{-6}$. For all the time-dependent datasets, the time-derivative time stepping scheme is used due to its better performance (see Figure 3(c)). We use the all2all training strategy [12] with a warm-up period. We first train the model with all the pairs with lead time $2\Delta t$ for a few epochs, then add the pairs with lead time $4\Delta t$ and train for more epochs, and continue until all lead times are covered. The warm-up period for the all2all training takes 20% of the training epochs.
We evaluate the estimates of the models against the ground-truth target functions following the metric used in [12]. We first normalize both the model estimate and the ground-truth function values by shifting and scaling them with global equation-wide mean and standard deviations. We then compute the relative $L^{1}$ error per each variable (group) as
$e_{k}^{m}={\frac{{\sum_{i}\left|u_{k,i}^{m}-\widehat{u}_{k,i}^{m}\right|}}{{\sum_{i}\left|u_{k,i}^{m}\right|}}},$ (34)
where ${u}_{k,i}^{m}$ and $\widehat{u}_{k,i}^{m}$ are respectively model estimate and ground-truth function values for sample $m$, variable group $k$, and spatial coordinate $i$. For all the Navier-Stokes and the Compressible Euler datasets, we also sum the components of the the velocity vector here. Hence, we get a scalar relative error $e_{k}^{m}$ per each target variable (group) and per each sample. We average over $k$ to get a scalar relative error per sample, and report the median error over all samples in the validation or the test dataset.
Appendix F Baseline architectures
F.1 CNO
A Convolutional Neural Operator (CNO) is a model that (approximately) maps bandlimited functions to bandlimited functions [37]. Let $\mathcal{B}_{w}$ be the space of bandlimited functions with the bandlimit $w$. A CNO is compositional mapping between function spaces $\mathcal{G}:\mathcal{B}_{w}(D)\to\mathcal{B}_{w}(D)$ and is defined as
$\mathcal{G}:u\mapsto P(u)=v_{0}\mapsto v_{1}\mapsto\ldots v_{L}\mapsto Q(v_{L})=\bar{u},$ (35)
where
$v_{l+1}=\mathcal{P}_{l}\circ\Sigma_{l}\circ\mathcal{K}_{l}(v_{l}),\quad 1\leq l\leq L-1,$ (36)
where $L$ is the number of CNO blocks and $D$ is the domain.
Here, $P$ is the lifting operator, $Q$ is the projection operator, $\mathcal{P}_{l}$ is either the upsampling or downsampling operator, $\mathcal{K}_{l}$ is the convolution operator and $\Sigma_{l}$ is the (modified) activation operator. For continuous time conditioning, we used lead time-conditioned instance normalization, defined by
$IN_{\alpha(t),\beta(t)}(\mathbf{v})(x)=\alpha(t)\odot IN(\mathbf{v})(x)+\beta(t)$ (37)
where $\mathbf{v}$ is an input function, $IN(\mathbf{v})$ is a regular instance normalization, and $\alpha(t)$ and $\beta(t)$ are (small) MLPs.
In all the benchmarks, the CNO architectures and training details utilized in this work are taken from the paper [12] and are as follows:
Trainable parameters: $39.1$M
Number of epochs: $400$
Lifting dimension: $54$
Up/downsampling layers: $4$
Residual blocks in the bottleneck: $6$
Residual blocks in the middle layers: $6$
Batch size: 32
Optimizer: AdamW
Weight decay: $10^{-6}$
Initial learning rate: $5\cdot 10^{-4}$
Scheduler: Linear with decreasing factor of $0.9$ every $10$ epochs
Early stopping: If the validation loss does not improve for 40 epochs
F.2 scOT
The scalable Operator Transformer (scOT) is a multiscale transformer model mapping discretized functions on a regular cartesian grid to discretized functions [12]. It patches the input functions and projects these visual tokens linearly into a latent space. A SwinV2 [24] encoder-decoder structure with ConvNeXt [25] residual blocks transforms the latent tokens in a hierarchical manner. SwinV2 transformer blocks apply cosine attention not on the full token grid, but only on windows of tokens, effectively reducing the computational burden of full attention. Downsampling operations in the encoder after each level merge four tokens into one, therefore coarsen the computational grid; upsampling in the decoder reverses these operations. After the decoder, the latent tokens are mapped back linearly into the physical space.
The model accepts continuous time or lead time inputs. It is conditioned through every single layer norm module by that lead time. The conditioned layer norm $\mathcal{N}$ applied to a latent token ${\bf v}\in\mathbb{R}^{C}$ at some level is implemented as follows:
$\mu({\bf v})$ $=\frac{1}{C}\sum_{l=1}^{C}({\bf v})_{l},\quad\sigma^{2}({\bf v})=\frac{1}{C}\sum_{l=1}^{C}(({\bf v})_{l}-\mu({\bf v}))^{2}$ (38)
$\mathcal{N}({\bf v},t)$ $=\alpha(t)\odot\frac{{\bf v}-\mu({\bf v})\cdot{\bf 1}_{C}}{\sigma^{2}({\bf v})}+\beta(t)$ (39)
with lead time $t\in\mathbb{R}_{\geq 0}$, ${\bf 1}_{C}\in\mathbb{R}^{C}$ a vector of ones, and $\alpha(t)$, $\beta(t)$ being learnable affine maps from lead time to gain and bias.
As with CNO, we utilize the model and training setup from [12]. For the sake of completeness, they are given as follows:
Trainable parameters: 40M
Patch size: 4
Embedding/latent dimension: 48
Window size: 16
Number of levels: 4
Number of SwinV2 blocks at each level: 8
Number of attention heads in each level: $[3,6,12,24]$
Number of ConvNeXt blocks in the residual connection: 2
Optimizer: AdamW
Scheduler: Cosine Decay with linear warmup of 20 epochs
Maximum learning rate: $5\cdot 10^{-4}$
Weight decay: $10^{-6}$
Batch size: 40
Number of epochs: $400$
Early stopping: If the validation loss does not improve for 40 epochs
Gradient clipping (maximal norm): 5
F.3 FNO
The Fourier neural operator (FNO) is a model that efficiently maps bandlimited functions to bandlimited functions via convolutions in Fourier space [20] and the fast Fourier transform (FFT). Similar to CNO, the FNO is a composition of lifting, convolution, and projection operators. The convolution operator, $\mathcal{K}_{l}(\phi)$, may be viewed as a multiplication of data, $v_{t}$, and a learnable kernel function, $R_{\phi}$, in Fourier space. Here, $\phi$ are learnable parameters of a neural network. The convolution operator may therefore be expressed as,
$\left(\mathcal{K}(\phi)v_{l}\right)(x)=\mathcal{F}^{-1}\left(R_{\phi}\cdot\left(\mathcal{F}v_{t}\right)\right)(x),$ (40)
where
$v_{l+1}(x)=(\sigma\circ\mathtt{IN})\left(Wv_{l}(x)+\left(\mathcal{K}(\phi)v_{t}\right)(x)\right),\quad 1\leq l\leq L-1,$ (41)
$W$ is a local integral operator and $\mathcal{F}$ is the Fourier transform, $\mathtt{IN}$ is a (lead time-conditioned) instance normalization. For efficiency, the Fourier transform is implemented via the fast Fourier transform (FFT) with a computational complexity of $O(n\log n)$, where $n$ is the number of points in the discretized mesh. Likewise, a low-pass filter removes high-frequency components of the Fourier transform, such that the weight tensor $R_{\phi}$ contains only $k_{max}<n$ modes. Because $k_{max}$ is independent of $n$, generally taken as a hyperparameter between $8$ and $32$, the inner multiplication has a complexity of $O(k_{max})$. As with CNO, we condition the model on lead time through instance normalizations.
The same setup as in [12] is used for this work:
Trainable parameters: 37M
Lifting dimension: 96
Number of Fourier layers: 5
Number of Fourier modes ($k_{max}$): 20
Optimizer: AdamW
Scheduler: Cosine Decay
Initial learning rate: $5\cdot 10^{-4}$
Weight decay: $10^{-6}$
Number of epochs: 400
Batch size: 40
Early stopping: If the validation loss does not improve for 40 epochs
F.4 Geo-FNO and FNO DSE
The FFT is limited to equispaced rectangular grids, and this restriction is inherited by the FNO. To apply the FNO to general geometries, the geometry-aware Fourier neural operator (Geo-FNO) [21] learns diffeomorphic deformations between general point clouds and such rectangular domains. Let $D_{a}$ be the physical domain, $D^{c}$ the computational domain, $x\in D_{a}$, and $\xi\in D^{c}$ be the corresponding mesh points, $\mathcal{T}_{a}=\{x^{(i)}\}\subset D_{a}$ the input meshes, $\mathcal{T}^{c}=\{\xi^{(i)}\}\subset D^{c}$ the computational meshes. The diffeomorphism $\phi_{a}$ transforms the points from the computational mesh to the physical mesh by
$\phi_{a}:D^{c}$ $\to D_{a},$ (42)
$\xi$ $\mapsto x,$
with inverse $\phi_{a}^{-1}$ and both $\phi_{a}$ and $\phi_{a}^{-1}$ are smooth. The Fourier transform is subsequently approximated as
$\left(\mathcal{F}_{a}v(k)\right)\approx\frac{1}{|\mathcal{T}_{a}|}\sum_{x\in\mathcal{T}_{a}}m(x)v(x)e^{2i\pi\langle\phi_{a}^{-1}(x),k\rangle},$ (43)
where $m(x)$ is a weight matrix based on the distribution of points from the input mesh.
In contrast, the FNO with discrete spectral evaluations (FNO DSE) [23] formulates a band-limited discrete Fourier transform based on the input mesh. Based on the observation that the number of Fourier modes employed by the FNO, $k_{max}$, is small and independent of the number of mesh points, $n$, a direct matrix-vector multiplication can be performed with $O(k_{max}n)$ complexity. For small $k_{max}$, this operation is more efficient than FFT. The discrete Fourier transform matrix for mesh points $\mathcal{T}_{a}=\{x^{(k)}\}\subset D_{a}$ is constructed by
$\mathbf{V}_{k,l}=\frac{1}{\sqrt{n}}\left[e^{-2\pi i\left({k}x^{(l)}\right)}\right]_{k,l=0}^{k_{max}-1,n-1}.$ (44)
Using this matrix as a basis for the bandlimited Fourier transform, the FNO DSE avoids the construction of a computational mesh. Instead, it works directly on the physical mesh.
In all the benchmarks, the Geo-FNO and FNO DSE architectures and model hyperparameters are taken from [23] and fixed for all experiments as follows:
Trainable parameters: $3.68$M (Geo-FNO), $3.30$ (FNO DSE)
Number of epochs: $1000$
Lifting dimension: $48$
Number of Fourier modes ($k_{max}$): $14$
Number of Fourier layers: $4$
Computational mesh size (Geo-FNO): $40\times 40$
Batch size: 16
Optimizer: Adam
Weight decay: $10^{-5}$
Initial learning rate: $5\cdot 10^{-4}$
Scheduler: Linear with decreasing factor of $0.9$ every $10$ epochs
Early stopping: If the validation loss does not improve for 40 epochs
Based on the works [23, 21], two sets of favorable training hyperparameters were tested for each experiment. The results of the best model are reported. The first set of hyperparameters is:
Initial learning rate: $1\cdot 10^{-3}$
Scheduler: Linear with decreasing factor of $0.5$ every $50$ epochs
The second set is:
Initial learning rate: $5\cdot 10^{-3}$
Scheduler: Linear with decreasing factor of $0.97$ every $10$ epochs
F.5 GINO
Geometry-Informed Neural Operator (GINO) [22] is composed of three main operators:
$\mathcal{G}\;=\;\mathcal{E}\,\circ\,\mathcal{P}\,\circ\,\mathcal{D},$ (45)
where $\mathcal{E}$ is the encoder, $\mathcal{P}$ is the processor, and $\mathcal{D}$ is the decoder. Specifically, $\mathcal{E}$ maps the input data from an unstructured mesh to a uniform latent grid using a graph neural operator (GNO). $\mathcal{P}$ processes the latent-grid features via the FNO. $\mathcal{D}$ maps the latent features back to the desired output mesh (or query points) using GNO.
GNO layers approximate an integral operator on a local neighborhood graph. Given a ball of radius $r>0$ around each point $x\in D$,
$v_{l}(x)\;=\;\int_{B_{r}(x)}\kappa\!\bigl{(}x,y\bigr{)}\,v_{l-1}(y)\,\mathrm{d}y,$ (46)
we discretize the integral by selecting $M$ neighbors, $\{y_{1},\dots,y_{M}\}\subset B_{r}(x)$, and form a Riemann sum:
$v_{l}(x)\;\approx\;\sum_{i=1}^{M}\kappa\!\bigl{(}x,y_{i}\bigr{)}\,v_{l-1}\!\bigl{(}y_{i}\bigr{)}\,\mu\!\bigl{(}y_{i}\bigr{)},$ (47)
where $\mu(\cdot)$ represents the local weights for integration. These localized integral updates are performed across a graph that connects each node $x$ to its neighbors in $B_{r}(x)$.
In all the benchmarks, the GINO architectures and model hyperparameters are taken from [22] and fixed for all experiments as follows:
Trainable parameters: $3.7$M
Number of epochs: $500$
Radius for GNO layers: $0.033$
GNO transform type: Linear
Projection channels: $256$
Lifting channels: $256$
FNO input channels: $64$
Number Fourier of modes: $16$
Number of Fourier layers: $3$
Optimizer: Adam
Batch size: $64$
Initial learning rate: $1\cdot 10^{-3}$
Scheduler: Linear decay with factor of $0.8$ every $100$ epochs
Appendix G Further experimental results
There are many hyperparameters that can affect the training stability as well as different aspects of the performance of a trained RIGNO. In the experiments presented in this work, we fix a set of architectural choices that perform fairly well for all the datasets and training sizes. These hyperparameters are listed in Table E.1. Note that these hyperparameters do not correspond to the best-performing architecture, but rather they guarantee a fair performance with all the desired properties. In Section G.2, we show that the performance can be improved by using larger models. In the experiments of this chapter, we perturb one or a few of these hyperparameters while leaving the rest fixed in order to study how the performance of the model changes with respect to the changed hyperparameters. The trainings are carried out using the settings described in Section E, unless otherwise stated.
G.1 Training Randomness
In order to quantify the dependence of the final results on training randomness, we train RIGNO five times, each time with a different seed for the pseudo-random number generator. We store the model with the lowest validation error during training as in all experiments, and check for fluctuations in the test error across all five repetitions of each experiment. In these experiments, we use the unstructured version of the datasets (random point clouds), we train using 512 training samples with batch size 8, and do not fine tune with fractional pairing. The results are summarized in Table G.1, where the maximum and minimum test errors across the repetitions are reported for each dataset. These fluctuations are attributed to the randomness in processes such as weight initializations, dataset shuffling, edge masking, and random sub-sampling of the coordinates. In the rest of the experiments, we only train each model once with a fixed training random seed, keeping in mind that the errors can slightly deviate because of training randomness.
Table G.1: Statistics of relative $L^{1}$ test errors with different random seeds. We train RIGNO-12 on 512 samples from each datasets. Each experiment is repeated 5 times.
Error [%]
Dataset Mean $\pm$ Standard deviation
NS-PwC 1.05 $\pm$ 0.040
NS-SVS 2.50 $\pm$ 0.157
G.2 Scaling with Model Size
In order to evaluate how the performance of RIGNO scales with model size, we start with RIGNO-12 as the base architecture and change the latent sizes (all at the same time), the number of message passing steps in the processor, the overlap factors (both at the same time), and the number of times that the regional mesh is sub-sampled when processor edges are built (number of regional edge levels). We train these models on 256 samples of the NS-PwC dataset without fractional pairing and calculate the error at $t_{14}$ on 128 test samples by using the AR-2 rollout scheme (defined in Section C.2). The results are plotted in Figure G.1.
We observe that the latent sizes and number of processor steps have the largest effect on the model capacity. The error consistently falls down when these hyperparameters are scaled, and we expect to get even lower errors with larger RIGNO architectures. With larger overlap factors, although the number of edges (and hence the computational cost) increases rapidly, the rate of convergence is relatively low. Therefore, the authors recommend using the minimum overlap, especially for the encoder. With more regional mesh levels, on the other hand, the number of edges only increases logarithmically, which is also reflected in the inference times. With 4 regional mesh levels, we obtain $3.75\%$ error, which is extremely lower than the $17.10\%$ error with only a single regional mesh level. The inference time, however, is only increased by $14\%$. These results show the effectiveness of longer edges in the construction of the regional mesh. However, depending on the space resolution of the inputs, there is saturation limit for this hyperparameter. If we consider a downsampling factor of 4, with $00016^{\prime}$ 000 input coordinates, for instance, we end up with a mesh with only $4$ nodes after 6 downsampling steps. With so few nodes we can only get 6 additional edges at most, which explains why the decreasing trend in the error stops at 4 downsampling steps.
Figure G.1: Test errors at $t=t_{14}$ with different strategies for scaling model size. The experiments are carried out on the NS-PwC dataset.
G.3 Temporal Continuity: Generalization to Unseen Lead Times
In order to assess the generalization of the trained models to unseen lead times $\tau$, we train RIGNO on the NS-PwC dataset with $2\Delta t$ time resolution, and test it on the test dataset with $\Delta t$ time resolution. With the incorporated mechanisms, namely time-conditioned normalization layers, all2all training, and fractional pairing fine-tuning, we expect to see a monotone increase in the error as we increase the lead time; which is very well captured in our tests as visualized in Figure G.2. Here, the lead time $\tau$ is fixed at each line in the plot, and the input time snapshot $t$ is changed. This explains why the first entries are missing for larger lead times. In the uppermost line ($\tau=10\Delta t$), for instance, the first entry corresponds to $t=0$, $\tau=10\Delta t$; the second entry to $t=1\Delta t$, $\tau=10\Delta t$; and so on.
In Figure G.2, the errors increase with larger lead times, which is expected since estimating the solution of the PDE with a higher lead time is inherently more difficult. In these experiments, we have removed the pairs with $\tau>8\Delta t$ from the all2all training pairs. Hence, $\tau=9\Delta t$ and $\tau=10\Delta t$ are considered as extrapolation in lead time. Since lead times $\tau=3\Delta t$, $\tau=5\Delta t$, and $\tau=7\Delta t$ are also not present in the all2all pairs, they are considered as interpolation in lead time. We can see that the monotonicity of the increase in the error is preserved for the extrapolated as well as the interpolated lead times. Moreover, the error increases in a controlled manner. The authors believe that this property is attributed to incorporation of time-conditioned normalization layers in the architecture and the use of time-derivative time-stepping strategy. However, with the smallest lead time $\tau=\Delta t$, it was observed that the monotonicity was not preserved without fine-tuning with the fractional pairing strategy.
Figure G.2: Direct test errors with different $\tau$ for the NS-PwC dataset. The vertical line indicates $t=t_{14}$, which is the last time snapshot seen by the model during training. Different lines denote different lead times. The gray lines correspond to lead times present in the training dataset; the green lines highlight interpolation in lead time; and the red lines indicate extrapolation in lead time. The model is trained on 512 samples with batch size 4.
G.4 Model Uncertainty
The edge masking technique, randomly disables a subset of the edges in each message-passing block. By adding randomness to the model, edge masking allows inferring the same model multiple times with the same input in order to generate an ensemble of estimates at a later time. The standard deviation of the generated ensemble can be interpreted as an indicator of model uncertainty on each output coordinate.
Figure G.3 visualizes the mean and the standard deviation of an ensemble of estimates at $t=t_{14}$ generated from input at $t=t_{0}$, along with the absolute error of the mean with respect to the ground-truth solution. More such visualizations are available in Section H. These visualizations show how the regions with the highest errors largely overlap with the regions where the ensemble has the highest deviations; which supports the interpretation of the deviations as a measure for model uncertainty, and shows that they are good indicators for inaccuracies in the estimates generated by the model.
Generally, the uncertainty of the model increases over autoregressive rollouts or with larger lead times. The model uncertainty can therefore be used to optimize the set of lead times in an autoregressive path. A possible algorithm would be to adaptively choose the largest lead time $\tau$ that results in uncertainties smaller than a threshold. Since model inference is fast and differentiable, grid-search algorithms can also be implemented. Similarly, the input discretization can adaptively be refined at each time step $t$ in the areas where the uncertainty of the model in its output is the highest.
Figure G.3: Mean and standard deviation of an ensemble of autoregressive estimates at $t_{14}$ for a test sample of the NS-SL dataset.
G.5 Robustness to Noise
In many applications, a learned operator needs to be inferred with inputs that are collected experimentally. Since measurement noise is inevitable in any experimental setup, robustness to noisy inputs is vital for such solution operators. Ideally, the operator should dampen the noise in the input; or at least the noise should not propagate in an uncontrolled way. To evaluate the robustness of RIGNO to input noise, we train it using clean, noise-free data, infer it with noisy inputs, and compute the error of the output with respect to clean ground-truth data. This setting emulates applications where the operator has been trained on noise-free high-fidelity simulation data and is applied on noisy experimental data.
(a) Heat-L-Sines
(b) NS-PwC
Figure G.4: Autoregressive test errors introduced by the noise in the input (relative). The y-axes show the difference in the relative error with noisy and clean inputs. The solid lines (best) show the results with edge masking and overlap factors 2.0. The dashed lines show the results with the same overlap factors but without edge masking during training. The dotted lines show the results with edge masking and overlap factors 1.0. All models are trained on 256 trajectories, without all2all training, and without fractional pairing fine-tuning.
For each input in the test dataset, we compute its standard deviation $\sigma_{\rm input}$, and add a Gaussian noise with standard deviation $\sigma_{\rm noise}=C_{noise}\sigma_{\rm input}$, with a noise level $C_{noise}\in\{0.005,0.02\}$. Starting with a noisy input, we apply the operators autoregressively with $\tau_{\rm max}=2\Delta t$. The errors in the resulting trajectories are partially due to the input noise, and partially to the approximation error of the operator. In order to distinguish between these two, we also unroll the operator with clean input ($C_{noise}=0$), and subtract the resulting errors from the errors with noisy inputs. This procedure gives a measure for the part of the error that is introduced by the noise in the input, denoted by noise-induced error. We present the relative noise induced error with respect to the approximation error: $|{e_{\rm noisy}-e_{\rm clean}}|/{e_{\rm clean}}$, where $e$ represents the autoregressive test error at $t=t_{14}$. The results are visualized in Figure G.4. With $2\%$ input noise at $t_{0}$, the noise is very well dampened to only $0.5\%$ at $t_{14}$ of the Heat-L-Sines dataset. With larger support sub-regions, the regional nodes aggregate data from a larger pool of physical nodes. Therefore, they are less sensitive to noisy inputs, and a stronger noise dampening is observed. Edge masking also plays an important role in noise dampening. When larger support sub-regions are used along with edge masking, $2\%$ input noise is dampened below $0.2\%$ at $t_{14}$.
Furthermore, noise dampening of RIGNO is not limited to Gaussian noise, as there has been no indication of this type of noise during training of the model. The authors believe that RIGNO is also robust to the errors that are introduced during unrolling; if we consider the introduced errors as noise with a certain distribution. In other words, RIGNO corrects itself within successive autoregressive steps. Therefore, for certain problems, small -but not too small- lead times might be advantageous to larger ones due to noise dampening.
Appendix H Visualizations
Visual inspection allows getting more insights about several aspects of the performance of RIGNO. Estimates and statistics of ensembles of estimates are visualized in this section for all datasets. All architecture hyperparameters and training details correspond to Section E. All the presented samples are from the test datasets, and have been obtained by applying the operator autoregressively with $\tau_{\rm max}=2\Delta t$ (where applicable).
Model estimates and their absolute errors, along with model input and the ground-truth output are visualized in Figures H.1-H.20. Visual inspection of model estimates for the CE-Gauss and CE-RP datasets reveals that there are two main reasons that explain the high errors for these datasets. First, model estimates are generally smoother than the ground-truth solutions. This inability to produce high-frequency estimates is reflected in the high errors where vortices are present. This is evident in the high-frequency details of velocity components in these model estimates. Second, although the model is able to capture shocks quite well at the right locations, it generates a smoother transition with smaller gradients. This is very well witnessed in the absolute estimate errors of Figure H.18, where the highest errors are located in small neighborhoods around shocks. The statistics of ensembles of estimates are visualized in Figures H.21-H.24. Estimate ensembles are generated by unrolling the operator randomly 20 times. Because of random edge masking, the estimates will slightly be different every time. The statistics of the estimate ensembles can act as indicators for model uncertainty. See Section G.4 for a broader discussion.
Figure H.1: Estimation of RIGNO-18 for a random test sample of the Heat-L-Sines dataset at $t_{14}$.
Figure H.2: Estimation of RIGNO-18 for a random test sample of the Wave-C-Sines dataset at $t_{14}$.
Figure H.3: Estimation of RIGNO-18 for a random test sample of the unstructured NS-Gauss dataset at $t_{14}$.
Figure H.4: Estimation of RIGNO-18 for a random test sample of the unstructured NS-PwC dataset at $t_{14}$.
Figure H.5: Estimation of RIGNO-18 for a random test sample of the unstructured NS-SL dataset at $t_{14}$.
Figure H.6: Estimation of RIGNO-18 for a random test sample of the unstructured NS-SVS dataset at $t_{14}$.
Figure H.7: Estimation of RIGNO-18 for a random test sample of the unstructured CE-Gauss dataset at $t_{14}$.
Figure H.8: Estimation of RIGNO-18 for a random test sample of the unstructured CE-RP dataset at $t_{14}$.
Figure H.9: Estimation of RIGNO-18 for a random test sample of the unstructured ACE dataset at $t_{14}$.
Figure H.10: Estimation of RIGNO-18 for a random test sample of the unstructured Wave-Layer dataset at $t_{14}$.
Figure H.11: Estimation of RIGNO-18 for a random test sample of the AF dataset.
Figure H.12: Estimation of RIGNO-18 for a random test sample of the Elasticity dataset.
Figure H.13: Estimation of RIGNO-18 for a random test sample of the NS-Gauss dataset at $t_{14}$.
Figure H.14: Estimation of RIGNO-18 for a random test sample of the NS-PwC dataset at $t_{14}$.
Figure H.15: Estimation of RIGNO-18 for a random test sample of the NS-SL dataset at $t_{14}$.
Figure H.16: Estimation of RIGNO-18 for a random test sample of the NS-SVS dataset at $t_{14}$.
Figure H.17: Estimation of RIGNO-18 for a random test sample of the CE-Gauss dataset at $t_{14}$.
Figure H.18: Estimation of RIGNO-18 for a random test sample of the CE-RP dataset at $t_{14}$.
Figure H.19: Estimation of RIGNO-18 for a random test sample of the ACE dataset at $t_{14}$.
Figure H.20: Estimation of RIGNO-18 for a random test sample of the Wave-Layer dataset at $t_{14}$.
Figure H.21: Model uncertainty with RIGNO-18 for a random test sample of the NS-Gauss dataset.
Figure H.22: Model uncertainty with RIGNO-18 for a random test sample of the AF dataset.
Figure H.23: Model uncertainty with RIGNO-18 for a random test sample of the CE-Gauss dataset.
Figure H.24: Model uncertainty with RIGNO-18 for a random test sample of the CE-RP dataset.